OpenStudy (anonymous):

A scientist counts 35 bacteria present in a culture and finds that the number of bacteria triples each hour. The function y = 35 • 3^x models the number of bacteria after x hours. Graph the function. (You do not need to submit the graph). Use the graph to estimate when there will be about 550 bacteria in the culture.

OpenStudy (anonymous):

550=35*3^x 550/35=3^x log(550/35)=xlog3 15.71428/log3=x x=2.5073178

OpenStudy (anonymous):

is that the answer?

OpenStudy (anonymous):

yes! put the x in the equation and you'll get 550.

OpenStudy (anonymous):

You don't even need to go through all that, if you do as you're supposed to the answer would be pretty clear. You can just see it and find the point where it looks like it is at. And there are places you can graph it online: http://www.quickmath.com/webMathematica3/quickmath/graphs/equations/basic.jsp#c=draw_basicgraphequation&v1=y%3D35+*+3%5Ex&v2=0&v3=4&v4=0&v5=600